can you help me find the points to graph thie function?
g(x)=1/3-4/3x

Answer by Alan3354(69443)

can you help me find the points to graph thie function?
g(x)=1/3-4/3x
-------------
sure.
g(x)=1/3-4/3x
g(x)=1/3-4x/3 = (1 - 4x)/3
------------------
Pick values of x that make the numerator multiples of 3 to make it easier.
eg, x = -2, 4, 10, etc
Solve for g(x)
Plot the points
Draw a line thru them.